A man walk 3 km in 1 hour.find the distance he can travel in 18 seconds.
Found 2 solutions by Alan3354, lenny460:
Answer by Alan3354(69443)

A man walk 3 km in 1 hour.find the distance he can travel in 18 seconds.
------------------
1 hour = 3600 seconds
----
d = 3 km/hr * 1 hr/3600 sec * 18 sec

Answer by lenny460(1073)

A man walk 3 km in 1 hour. Find the distance he can travel in 18 seconds ?

1 km = 1000 meters
3km = 3000 meters
1 hour = 60 minutes
1 minute = 60 seconds
Therefore:
1 hour = 60 * 60 = 3600 seconds

3000 / 3600 = 0.83 meters per second
He walks 0.83 meters per second

In 18 seconds:
0.83 * 18 = 15 meters

Answer:
In 18 seconds he will walk 15 meters.









Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com